OpenStudy (abdullah1995):

A forece acts at the origin of a coordinate system in a direction defined by the angles theta(x) = 75, Theta(z) = 130. Knowing that the y component of the force is +300 lb, determine, (a) the angle Theta(y) (b) the other components and the magnitude of the force.

OpenStudy (abdullah1995):

@ganeshie8

OpenStudy (irishboy123):

you have \(\vec F = F_T<cos \theta_x,cos \theta_y,cos \theta_z>\) and \(\vec F \bullet \vec F = F^2_T <cos \theta_x,cos \theta_y,cos \theta_z> \bullet <cos \theta_x,cos \theta_y,cos \theta_z> = F^2_T\)

OpenStudy (ybarrap):

You can get \(F_T\) from the information in the diagram below  Then the other components can be calculated: $$ F_x=F_T\cos 75\\ F_z=F_T\cos 130\\ $$ Does this make sense?